# Quanta Diffusion

### v1

@@ 3.1 Single Photon Detector Forward Model
Given the quanta exposure, $\mathbf{I}^{\textbf{GT}}$ that is dependent on the photon flux and the exposure time, the observed signal by the sensor can be modeled as a Poisson-Gaussian random variable where the Poisson represents the photon arrival process and the Gaussian models the read noise. The read-out process involves various sorts of distortions and the Analog-to-Digital Converter (ADC) to represent the overall read-out as integers between $[0,\text{N}]$ where $\text{N}=2^{\text{Nbits}}-1$, Nbits indicating the bit-depth of the image sensor. The final sensor readout, Y, can be mathematically represented as,
@@ 4 Quanta Diffusion
Building on the discussed challenges and solutions, we propose a multi-frame physics-based generative framework, Quanta Diffusion (QuDi), illustrated in Figure 2. Inspired by QUIVER [2] and classical quanta burst methods [1, 19], QuDi leverages the core principles of these approaches. Figure 2 outlines a single step of the reverse diffusion process. The conditional reverse diffusion begins with Gaussian noise ($\mathbf{x}_{\text{T}}\sim\mathcal{N}(0,\mathbf{I})$) and noisy quanta frames ($\mathbf{y}\sim\text{Eq.}$. 1 $\in\mathds{R}^{\text{H}\times\text{W}\times\text{N}}$, where N is the number of frames) to generate a restored noiseless frame $\mathbf{\hat{x}}_{\theta}(\mathbf{x}_{\text{T}},\mathbf{y})$. At any arbitrary iteration, the QuDi architecture can be divided into $4$ main stages as described below:
Predenoising to increase SNR. Existing classical quanta burst methods like [1] utilize naive averaging method to increase the Signal-to-Noise Ratio (SNR). Although naive averaging proves to be the best method when the motion between the frames is negligible, it produces significant motion blur in the case of fast moving scenes, in turn affecting the flow estimation accuracy. Following QUIVER’s analogy [2], we design a computationally efficient single image denoiser as it helps remove dependency on inter-frame motion. As shown in Figure 2, though the initial denoised frames look heavily oversmoothened, it provides extensive details on the overall structural view of the scene, thus helping to reasonably estimate the optical flow. The single image denoiser is built upon the residual dense blocks [20] due to its simple yet effective design. While we denoise the SPD captured frames, we also perform $3$D shallow-feature extraction on the Gaussian noise induced high-bit frames $\mathbf{x}_{\text{t}}$ to extract useful spatio-temporal details. We adopt DDPM’s self-conditioning mechanism [14] to ensure stable iterative generation.
@@
Loss Function. We train QuDi with the conventional L2222 loss function with total variation and perceptual loss as regularizers. Mathematically, the overall loss function for training can be represented as,
@@
where $\mathbf{x}^{\text{r}}_{\text{t}}$, $\mathbf{x}^{\text{d}}_{\text{t}}$ are the restored and predenoised frames, respectively. Lastly, $\mathcal{L}$ is a combination of L2222 and total variation losses.
@@ 5.1 Experimental Settings
Baselines. We compare the proposed method with eight existing dynamic scene reconstruction algorithms, namely Transform Denoise [5], QBP [1], Student-Teacher [4], RVRT [10], EMVD [22], FloRNN [9], MemDeblur [23], Spk2ImgNet [24], and QUIVER [2]. We also add an off-the-shelf denoiser BM3333D [25] to QBP, denoted QBP (+BM3333D), as a baseline for comparison. As we will discuss in Section 5.2, QuDi beats all the baselines, both quantitatively and qualitatively.
Training. We adopt the $2000$ fps high-speed dataset, I2222-$2000$fps open-sourced by [2] and employ it as the training dataset for all the deep-learning models in our experiments. Each training sample is fetched on the fly from each clip. A training sample here is defined as a tuple containing the ground-truth/target frames and the $3$-bit quanta frames simulated at $3.25,9.75,26$ photons-per-pixel (PPP) ($\sim 1$ lux assuming a $1.1\mu\text{m}^{2}$ pixel pitch and a $1/2000$ second exposure time) using the image formation model described in Section 3.1.
We utilize the function mentioned in Eq. 4 as the cost function for training QuDi with regularization parameters $\lambda_{1}=1$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.3$. The training data is extracted with a patch size of $256\times 256$ and a batch size of $32$. The network is trained using Adam optimizer along with a cosine scheduled learning rate of $2\times 10^{-5}$. QuDi takes approximately $3$ days to train on $8$ NVIDIA A100100100100 GPUs using Pytorch.
Testing. To effectively analyze the performance of various methods, we utilize $31$ videos from I2222-$2000$FPS containing various motion types, shapes, and speeds. Lastly, to measure the performance on real-world data, we collect and utilize the real-SPAD data using a SPAD sensor [26]. More details will be discussed in Section 5.2. For QuDi, with $11$ SPD frames as input, generating $11$ output frames over $10$ diffusion time-steps takes approximately $2.5$ seconds on an A100 GPU.
@@ 5.2 Results
Table 1: Performance Comparison on the I2222-$2000$fps dataset. QuDi out-performs all the baselines by a significant margin.
Fig. 3: Synthetic Data Results. Visual comparisons of reconstructed frames using test videos from the I2222-$2000$FPS dataset. For a fair comparison, all methods use $11$ $3$-bit quanta frames simulated at 3.25 PPP per frame ($\sim 1$ Lux) to generate a restored frame. Best viewed in zoom.

### v2

@@ 3.1 Single Photon Detector Forward Model
Given the quanta exposure, $\mathbf{I}^{\textbf{GT}}$ that is dependent on the photon flux and the exposure time, the observed signal by the sensor can be modeled as a Poisson-Gaussian random variable where the Poisson represents the photon arrival process and the Gaussian models the read noise. The read-out process involves various sorts of distortions and the Analog-to-Digital Converter (ADC) to represent the overall read-out as integers between $0,\text{N}]$ where $\text{N}=2^{\text{Nbits}}-1$, Nbits indicating the bit-depth of the image sensor. The final sensor readout, Y, can be mathematically represented as,
@@ 4 Quanta Diffusion
Building on the discussed challenges and solutions, we propose a multi-frame physics-based generative framework, Quanta Diffusion (QuDi), illustrated in Figure 2. Inspired by QUIVER [2] and classical quanta burst methods [1, 19], QuDi leverages the core principles of these approaches. Figure 2 outlines a single step of the reverse diffusion process. The conditional reverse diffusion begins with Gaussian noise ($\mathbf{x}_{\text{T}}\sim\mathcal{N}(0,\mathbf{I})$) and noisy quanta frames ($\mathbf{y}\sim\text{Eq.}$ 1 $\in\mathds{R}^{\text{H}\times\text{W}\times\text{N}}$, where N is the number of frames) to generate a restored noiseless frame $\mathbf{\hat{x}}_{\theta}(\mathbf{x}_{\text{T}},\mathbf{y})$. At any arbitrary iteration, the QuDi architecture can be divided into $4$ main stages as described below:
Predenoising to increase SNR. Existing classical quanta burst methods like [1] utilize naive averaging method to increase the Signal-to-Noise Ratio (SNR). Although naive averaging proves to be the best method when the motion between the frames is negligible, it produces significant motion blur in the case of fast moving scenes, in turn affecting the flow estimation accuracy. Following QUIVER’s analogy [2], we design a computationally efficient single image denoiser as it helps remove dependency on inter-frame motion. As shown in Figure 2, though the initial denoised frames look heavily oversmoothened, it provides extensive details on the overall structural view of the scene, thus helping to reasonably estimate the optical flow. The single image denoiser is built upon the residual dense blocks [20] due to its simple yet effective design. While we denoise the SPD captured frames, we also perform 33D shallow-feature extraction on the Gaussian noise induced high-bit frames $\mathbf{x}_{\text{t}}$ to extract useful spatio-temporal details. We adopt DDPM’s self-conditioning mechanism [14] to ensure stable iterative generation.
@@
Loss Function. We train QuDi with the conventional L22 loss function with total variation and perceptual loss as regularizers. Mathematically, the overall loss function for training can be represented as,
@@
where $\mathbf{x}^{\text{r}}_{\text{t}}$, $\mathbf{x}^{\text{d}}_{\text{t}}$ are the restored and predenoised frames, respectively. Lastly, $\mathcal{L}$ is a combination of L22 and total variation losses.
@@ 5.1 Experimental Settings
Baselines. We compare the proposed method with eight existing dynamic scene reconstruction algorithms, namely Transform Denoise [5], QBP [1], Student-Teacher [4], RVRT [10], EMVD [22], FloRNN [9], MemDeblur [23], Spk2ImgNet [24], and QUIVER [2]. We also add an off-the-shelf denoiser BM33D [25] to QBP, denoted QBP (+BM33D), as a baseline for comparison. As we will discuss in Section 5.2, QuDi beats all the baselines, both quantitatively and qualitatively.
Training. We adopt the $2000$ fps high-speed dataset, I22-20002000fps open-sourced by [2] and employ it as the training dataset for all the deep-learning models in our experiments. Each training sample is fetched on the fly from each clip. A training sample here is defined as a tuple containing the ground-truth/target frames and the $3$-bit quanta frames simulated at $3.25,9.75,26$ photons-per-pixel (PPP) ($\sim 1$ lux assuming a $1.1\mu\text{m}^{2}$ pixel pitch and a $1/2000$ second exposure time) using the image formation model described in Section 3.1.
We utilize the function mentioned in Eq. 4 as the cost function for training QuDi with regularization parameters $\lambda_{1}=1$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.3$. The training data is extracted with a patch size of $256\times 256$ and a batch size of $32$. The network is trained using Adam optimizer along with a cosine scheduled learning rate of $2\times 10^{-5}$. QuDi takes approximately $3$ days to train on $8$ NVIDIA A100100 GPUs using Pytorch.
Testing. To effectively analyze the performance of various methods, we utilize $31$ videos from I22-20002000FPS containing various motion types, shapes, and speeds. Lastly, to measure the performance on real-world data, we collect and utilize the real-SPAD data using a SPAD sensor [26]. More details will be discussed in Section 5.2. For QuDi, with $11$ SPD frames as input, generating $11$ output frames over $10$ diffusion time-steps takes approximately $2.5$ seconds on an A100 GPU.
@@ 5.2 Results
Table 1: Performance Comparison on the I22-20002000fps dataset. QuDi out-performs all the baselines by a significant margin.
Fig. 3: Synthetic Data Results. Visual comparisons of reconstructed frames using test videos from the I22-20002000FPS dataset. For a fair comparison, all methods use $11$ 33-bit quanta frames simulated at 3.25 PPP per frame ($\sim 1$ Lux) to generate a restored frame. Best viewed in zoom.